OpenStudy (anonymous):

A package of modeling clay was shaped like a cone with a radius of 24 cm and a height of 6 cm. Miriam used all the clay to make a cylinder with a radius of 16 cm. What was the height of the cylinder? Use 3.14 to approximate pi, and express your final answer IN TENTHS. __cm

OpenStudy (anonymous):

@dan815

OpenStudy (welshfella):

here you have to realise that the volume of the cone will be same as the volume of the cylinder

OpenStudy (anonymous):

so the volume is 3619.11 so i would divide that by 16^2 and 3.14?

OpenStudy (welshfella):

Volume of the cone = (1/3) pi r^2 h = (1/3) pi 24^2 * 6 volume of the cylinder is pi r^2 h = pi 16^2 h equate these 2 and solve for h

OpenStudy (anonymous):

Is the height 4.5 cm?

OpenStudy (welshfella):

yes

OpenStudy (anonymous):

Thanks I didn't know that the cylinder and Cone had the same volume!

OpenStudy (welshfella):

they must have because they are both made of the same amount of clay

OpenStudy (anonymous):

Yeah that makes sense i wasn't reading like that! Thanks again!